OpenStudy (el_arrow):

find the area of the surface generated by revolving the curve on the interval 0

OpenStudy (el_arrow):

x = t and y= 7t

OpenStudy (el_arrow):

@perl help me please

OpenStudy (perl):

what is the given axis

OpenStudy (el_arrow):

it doesnt tell me

OpenStudy (el_arrow):

just says about the given axis....

OpenStudy (el_arrow):

when that happens can i use either axis y or x?

OpenStudy (lilshane):

your good @perl

OpenStudy (el_arrow):

yeah he is one of the best

OpenStudy (perl):

we might get difference answers if you rotate about the y axis versus rotating about x axis

OpenStudy (el_arrow):

my teacher gave us this problem i guess he forgot to pick an axis

OpenStudy (lilshane):

per i wish i cod be mod or ambassdor

OpenStudy (perl):

$$\Large{ \text{rotation about x axis: }\int2\pi y ds :\\ \text{rotation about y axis: }\int 2 \pi x ds : } $$

OpenStudy (el_arrow):

okay and i'll get the same answer from both of them

OpenStudy (perl):

$$\Large{ \text{rotation about x axis: }\\ \int_{0}^{3}2\pi y \sqrt{ \left( \frac{dx}{dt} \right)^2+ \left( \frac{dy}{dt} \right) ^2} ~dt \\ } $$

OpenStudy (perl):

$$\Large{x = t ,y= 7t\\ \text{rotation about x axis: }\\ \int_{t_1}^{t_2}2\pi y \sqrt{ \left( \frac{dx}{dt} \right)^2+ \left( \frac{dy}{dt} \right) ^2} ~dt \\ =\int_{0}^{3}2\pi \cdot 7t \sqrt{ \left( 1 \right)^2+ \left( 7 \right) ^2} ~dt \\ \\ } $$

OpenStudy (el_arrow):

dx = 1 and dy = 7 right?

OpenStudy (el_arrow):

okay and i dont need to change the integrals for this problem correct?

OpenStudy (el_arrow):

the t2 and t1 remain 3 and 0 right

OpenStudy (perl):

right

OpenStudy (perl):

surface of area of revolution does change

OpenStudy (el_arrow):

what do you mean

OpenStudy (perl):

$$ \Large{x = t ,y= 7t\\ \text{rotation about x axis: }\\ \int_{t_1}^{t_2}2\pi y \sqrt{ \left( \frac{dx}{dt} \right)^2+ \left( \frac{dy}{dt} \right) ^2} ~dt \\ =\int_{0}^{3}2\pi \cdot 7t \sqrt{ \left( 1 \right)^2+ \left( 7 \right) ^2} ~dt \\ =\sqrt{50}\int_{0}^{3}2\pi \cdot 7t ~dt \\ =315 \pi \sqrt 2 \\ \therefore \\ \text{rotation about y axis: }\\ \int_{t_1}^{t_2}2\pi x \sqrt{ \left( \frac{dx}{dt} \right)^2+ \left( \frac{dy}{dt} \right) ^2} ~dt \\ =\int_{0}^{3}2\pi \cdot t \sqrt{ \left( 1 \right)^2+ \left( 7 \right) ^2} ~dt \\ =\sqrt{50}\int_{0}^{3}2\pi \cdot t ~dt \\ =45 \pi \sqrt 2 \\ } $$

OpenStudy (el_arrow):

okay so the answers are not the same

OpenStudy (el_arrow):

if i ask one more question on here would you help me it'll be the last one for the night

OpenStudy (el_arrow):

i just want to know if my work is right

OpenStudy (el_arrow):

could you?

OpenStudy (perl):

ok

OpenStudy (el_arrow):

alright thank you

OpenStudy (el_arrow):

i am going to draw it out

OpenStudy (el_arrow):

this is problem find the arc length of the curve on the interval 0<t<6

OpenStudy (el_arrow):

x = sqrt(t) and y = 9t-6

OpenStudy (el_arrow):

for dx/dt = 1/2sqrt(t) and dy/dt = 9

OpenStudy (el_arrow):